The school book store sold 8 more pencils than pens one day. The cost of a pencil is $.05, and the cost of a pen is $.20. If the day's sales of pens and pencils totaled $8.90, how many pencils were sold?
Answer by Nate(3500)

amount of pens sold = p
amount of pencils sold = p + 8
0.05(p + 8) + 0.20(p) = 8.90 each pen is $0.20 while each pencil is $0.05
0.05p + 0.4 + 0.20p = 8.90
0.25p = 8.50
amount of pens: 34
amount of pencils: 34 + 8 = 42